# Revisiting Privacy, Utility, and Efficiency Trade-offs when Fine-Tuning Large Language Models

### v1

@@ II. Examining memorized sequences from [3]:
In the right chart in Figure 4, the memorized strings are predicted mostly from GitHub, followed by ArXiv, while the rest of the sources are largely under-represented. Herein, both GitHub and ArXiv are relatively in the middle range in terms of contents in the original dataset in the top chart. However, analyzing the typical data in these sections, GitHub appears as a source of structured format code with repeated predictable patterns, which is commonly tagged as non-sensitive data. Similarly, the Pile dataset includes files uploaded to ArXiv, since is a common typesetting language for scientific research papers [13]. As such, highly memorized strings in the Pile dataset are non-sensitive in nature.
Validating GPT-4 predictions. GPT-4 predictions may be erroneous. Hence, we conduct a verification test to evaluate the accuracy of GPT-4’s predictions. For this assessment, we sample $200$ random strings from each of the 22 sections of the Pile dataset [13], and prompt GPT-4 to predict the source of the strings. Unlike the previous experiment, the ground-truth of string source is known in this validation experiment. Figure 5 illustrates the accuracy for each section, indicating that $50\%$ of the sections exhibit an accuracy rate of at least $90\%$ with $4.5\%$ being the base accuracy of a random predictor. Furthermore, GPT-4 predicts the correct source on an average of $78\%$ strings across all $22$ sections of the Pile dataset. In addition, misclassified strings are often assigned to sections of a similar category, e.g., NIH Explorer misclassified as PubMed Central (more details in Appendix 10). Therefore, the GPT-4 predictions can be considered as reliable.
@@
Figure 6: Most participants classified the memorized sequences detected by [3] as non-sensitive, with fewer than 10% marking them as privacy-sensitive, indicating that the perceived privacy risk of these strings is generally low.
Finally, we conduct a human survey on Prolific to evaluate the extent of sensitive information present in randomly chosen $100$ memorized strings from [3]. The survey results are summarized in Figure 6. The majority of participants classified memorized strings as non-sensitive, while $<10\%$ participants disagree and mark the strings as containing privacy-sensitive information. Therefore, most crowdsourced participants do not perceive the sampled stings as containing privacy-sensitive content. Further details on the setup and results are provided in Appendix 12. Thus, by distinguishing between sensitive and non-sensitive entities, we demonstrate a deeper understanding of actual privacy threat.
4 Privacy-Utility-Efficiency tradeoffs in different fine-tuning methods
@@
We use the distinction between sensitive and non-sensitive tokens to study the privacy and utility impact of training models with three different fine-tuning methods: full fine-tuning (FFT), Differentially Privacy (DP), and Low-Rank Adaptation (LoRA). We also investigate the computational efficiency of each method. Our goal is to answer the following questions: “How prone is each method to recollecting the sensitive parts of the training data? (Privacy)” , “How effective is each method at predicting non-sensitive parts of test data? (Utility)” , “What is the computational cost associated with each method? (Efficiency)”.
@@
To answer these questions, we use each fine-tuning method to train three models from different families, Pythia-1B [4], Gemma-2B [45] and Llama2-7B [47], on two datasets, CustomerSim and SynBio. More information about the datasets and our methodology for distinguishing between sensitive and non-sensitive tokens can be found in Section 3.2. We train the models for 50 epochs on each dataset. Details on hyperparameters can be found in Appendix 9.
@@
For each fine-tuning method, model and dataset we report three metrics: 1) privacy as the loss on sensitive tokens (annotated by GPT-4) in the training data, 2) utility as the loss on non-sensitive tokens (the remaining tokens) on a held-out test set, and 3) efficiency based on the relative amount of computation and memory usage of each method. Additionally, to assess how fine-tuning affects the underlying abilities and knowledge of the base model, we measure the performance of the fine-tuned Gemma model (trained with CustomerSim data) on general language understanding benchmarks: SCIQ [21], a dataset of over 13,000 crowdsourced questions on Physics, Chemistry, and Biology; MMLU [17, 16], a large multi-task dataset covering various domains of knowledge; and Hellaswag [53], a dataset for commonsense inference.
Update rules: For each fine-tuning method, we describe how updated weights $W_{t+1}$ are computed from the previous weights $W_{t}$ in each step, where $W_{0}$ are the weights of the pre-trained base model before fine-tuning. We use $X$ to refer to a batch of $|X|$ datapoints and $x_{i}$ to refer to individual datapoint, $\mathcal{M}_{W}$ to refer to the model parameterized by weights $W$, and $\mathcal{L}(\mathcal{M}_{W}(X),X)$ to refer to the autoregressive cross-entropy loss of the model on data $X$. We denote by $\nabla_{W}\mathcal{L}(...)$ as the gradient of the loss wrt. weights $W$ and $\eta$ is the learning rate.
@@
Efficiency: The efficiency of each method is determined by the amount of computation it requires, and also other factors such as memory requirements, which can affect the usable batch-size and thus the overall training throughput. Following [23], we estimate the amount of training compute ($C$) in floating point operations (FLOPs) for full fine-tuning as $C_{\text{FFT}}=6DN$, where $D$ is the number of training tokens and $N$, the number of model parameters. For each method, we report its compute requirements relative to the FFT-baseline based on measurements using the PyTorch profiler [^3] https://pytorch.org/docs/stable/profiler. We also comment on other factors that affect training throughput.
4.1 Method 1 : Full fine-tuning
@@
Efficiency: FFT serves as our efficiency baseline. It has moderate compute requirements (discussed above), and relatively high memory requirements, since in addition to the input-dependent activations, we need to keep four numbers per model parameter in GPU memory: the parameter value, its gradient, and two optimizer states (first and second moments of the gradient for Adam [26]).
@@
(a) Pythia (CustomerSim)
(b) Gemma (CustomerSim)
(c) Llama2 (CustomerSim)
(d) Pythia (SynBio)
(e) Gemma (SynBio)
(f) Llama2 (SynBio)
Figure 9: Privacy-utility trade-offs for DP across models and datasets. Top row: For the CustomerSim dataset, DP enhances privacy across all models, though higher noise scales reduce utility, especially for larger models. Bottom row: For the SynBio dataset, DP similarly improves privacy, with utility declines at higher noise levels. Gemma shows a unique utility drop after two epochs, highlighting the complexity of privacy-utility trade-off.
4.2 Method 2 : Differential Privacy
Differential Privacy (DP) algorithms  [1] aim to safeguard the privacy of individual training data points by limiting their influence on the gradient updates during training.
Update rules: DP clips the $l_{2}$ norm of each data point’s gradient at a threshold $T$, followed by adding noise with magnitude $\sigma$ to each clipped gradient. The purpose of clipping is to reduce data sensitivity by ensuring that the impact of data points with high gradient magnitudes on the model parameters is limited. Adding noise further obscures the contribution of individual data points, making it difficult to infer specific data points from the model.
@@ 4.2 Method 2 : Differential Privacy
We vary the noise hyperparameter $\sigma$ for the experiments. The clipping gradient norm $T$ is fixed at $10^{-2}$, as in [44].
@@
Figure 10: Fine-tuning Gemma with DP results in a substantial accuracy declines of 70%, 10%, and 30% across the benchmarks.
Privacy-Utility trade-off: Figures 9(a), 9(b), and 9(c) illustrate the privacy-utility trade-off when varying the noise $\sigma$ on the CustomerSim dataset across the Pythia, Gemma, and Llama2 models. As in previous plots, each curve begins with the performance of the pre-trained model. It is evident that DP maintains privacy effectively with minimal degradation. But there is a trade-off between privacy and utility. For all models, lower noise values such as $\sigma=0.1$ are able to achieve better utility than higher ones such as $\sigma=0.5$ and $\sigma=0.9$, but they also decrease privacy more. While the total amount of utility achievable with DP is limited, especially for the larger Llama2-7B model, overall, it provides good privacy-utility trade-offs. Similar trends can be observed for the SynBio dataset in Figures 9(d) and 9(f), but for Figure 9(e) (with Gemma) utility declines after approximately two epochs.
Impact on benchmark datasets: Figures 10(a), 10(b) and 10(c) show the benchmark accuracy over epochs for the DP fine-tuned Gemma model with $\sigma=0.1$ on SCIQ, MMLU, and HellaSwag datasets, respectively. Accuracy drops gradually with increasing epochs for all benchmarks, stabilizing at lower levels, indicating that fine-tuning with DP significantly reduces the knowledge retention capacity.
Efficiency: Differential privacy comes with a high computational cost, since the gradients of each datapoint need norm and clipping computations, and additional noise values are added. Empirically, we observe a relative FLOPs requirement of $C_{\text{DP}}/C_{\text{FFT}}=1.33$. Additionally, the per-sample operations required for clipping mean that we need to keep copies of the gradient for each datapoint in memory, which requires substantially more GPU memory, which decreases the feasible batch size and the overall training throughput even further.
@@
(d) Pythia (Rank 32)
(e) Gemma (Rank 32)
(f) Llama2 (Rank 32)
Figure 11: Privacy-utility trade-offs for LoRA fine-tuning on the CustomerSim dataset. Top row: LoRA with rank 16 shows that smaller models achieve a better privacy-utility trade-off, while larger models retain utility but experience reduced privacy with more epochs. Bottom row: LoRA with rank 32 yields similar results, with smaller models performing better in privacy-utility trade-off and larger models maintaining utility but with privacy reduction as epochs increase.
Takeaway: DP offers a reasonable privacy-utility trade-off and is less susceptible to learning sensitive data. Increasing noise rates lead to poor utility across all models. Convergence with DP is not guaranteed, particularly in larger models such as Llama2. The privacy gains come at the cost of efficiency. Additionally, fine-tuning with DP leads to quick decline in the benchmark performance. The degree of measures along the tradeoff, knowledge retention, and efficiency are:
Utility-privacy trade-offs: moderate
Retention of base performance: poor
Efficiency: poor
(a) Pythia (Rank 16)
(b) Gemma (Rank 16)
(c) Llama2 (Rank 16)
(d) Pythia (Rank 32)
(e) Gemma (Rank 32)
(f) Llama2 (Rank 32)
Figure 12: Privacy-utility trade-offs for LoRA fine-tuning on the SynBio dataset. Top row: LoRA with rank 16 shows that smaller models achieve a better privacy-utility trade-off, while larger models retain utility but experience reduced privacy with more epochs. Bottom row: LoRA with rank 32 yields similar results, with smaller models performing better in privacy-utility trade-off and larger models maintaining utility but with privacy reduction as epochs increase.
@@
4.3 Method 3 : Low-Rank Adaptation (LoRA)
LoRA [20] is a parameter-efficient fine-tuning method developed to reduce the compute and memory requirements of fine-tuning LLMs, and to reduce the size of storing fine-tuned checkpoints. It enjoys large popularity for LLM fine-tuning.
@@ 4.3 Method 3 : Low-Rank Adaptation (LoRA)
We hypothesize that LoRA’s low-rank updates restrict the model’s capacity to memorize precise details, which could have an effect similar to the noisy updates in DP. Therefore, LoRA has the potential to provide better privacy-utility trade-offs than FFT, similar to DP, while also being computationally more efficient.
Privacy-Utility trade-off: We investigate the effects of varying both the rank $r$ and scaling parameter $\alpha$ of LoRA. We use common rank values of $16$ and $32$ and $\alpha\in\{16,32,64,128\}$. While LoRA has been explored for privacy in conjunction with DP [51], there has been no prior work that specifically examines the privacy benefits of LoRA alone.
Figures 11(a)11(c) show the trade-off with rank $16$ and varying parameters of $\alpha$ for the CustomerSim dataset across Pythia, Gemma and Llama2 models. As seen in Figures 11(a) and 11(b), smaller-scale models (Pythia and Gemma) exhibit a better privacy-utility trade-off when the rank and $\alpha$ values are equal, compared to the other configurations. For the larger model, the privacy declines at later epochs, while utility is mostly retained. This trend is similar to the one observed for rank $32$ in Figures 11(d)11(f). We also analyze the trade-off for the SynBio dataset in Figures 12(a)12(c) for rank 16, and Figures 12(d)12(f) for rank 32, which make similar observations. However, due to the more unstructured nature of the SynBio dataset, there is a larger reduction in utility after certain epochs compared to the CustomerSim dataset.
@@
Efficiency: During training, LoRA has slightly larger compute requirements for the forward pass than full fine-tuning, since additional FLOPs are required for the adapter matrices $\Delta W$, though they are much smaller than the full base weight $W_{0}$. However, during the backward pass, LoRA requires less compute, since no gradients need to be computed for the base weights $W_{0}$. We observe a relative FLOPs requirement of $C_{\text{LoRA}}/C_{\text{FFT}}\approx 0.65$. The original paper [20] reports a 25% speedup during training. LoRA has needs of less GPU memory than FFT, since no optimizer states and gradients need to be stored for the base weights $W_{0}$, which makes it possible to run it with larger batch-sizes and thus an overall increased training throughput.
(a) Pythia (CustomerSim)
(b) Gemma (CustomerSim)
(c) Llama2 (CustomerSim)
(d) Pythia (SynBio)
(e) Gemma (SynBio)
(f) Llama2 (SynBio)
Figure 14: Privacy-utility tradeoffs for different fine-tuning methods on the CustomerSim and SynBio datasets. Top row: On the CustomerSim dataset, DP achieves high privacy with a significant computational cost, full fine-tuning maximizes utility but offers less privacy, and LoRA provides a balanced trade-off with the lowest computational overhead. Bottom row: Similar trends are observed for the SynBio dataset.
Takeaway: We are one of the first works to explore the privacy benefits of parameter-efficient fine-tuning methods, particularly LoRA. We vary LoRA’s $r$ and $\alpha$ hyperparameters and observe that all configurations are able to achieve high utility, while especially lower $r$ and $\alpha$ values also preserve a high degree of privacy. For each model, the optimal privacy-utility trade-off value is achieved with $r=\alpha$. In addition, we observe that LoRA, after being fine-tuned on the CustomerSim dataset, did not lose much of its abilities on the benchmark datasets, maintaining results comparable to those of pre-trained model. Finally, LoRA is much more computationally and memory efficient than FFT and especially DP. Overall it provides the best trade-offs in terms of utility, privacy and efficiency and shows that privacy can be achieved without additional computational costs. The degree of measures along the trade-off, knowledge retention, and efficiency are:

### v2

@@ II. Examining memorized sequences from [biderman2024emergent]:
In the right chart in Figure 4, the memorized strings are predicted mostly from GitHub, followed by ArXiv, while the rest of the sources are largely under-represented. Herein, both GitHub and ArXiv are relatively in the middle range in terms of contents in the original dataset in the top chart. However, analyzing the typical data in these sections, GitHub appears as a source of structured format code with repeated predictable patterns, which is commonly tagged as non-sensitive data. Similarly, the Pile dataset includes files uploaded to ArXiv, since is a common typesetting language for scientific research papers [gao2020pile]. As such, highly memorized strings in the Pile dataset are non-sensitive in nature.
Validating GPT-4 predictions. GPT-4 predictions may be erroneous. Hence, we conduct a verification test to evaluate the accuracy of GPT-4’s predictions. For this assessment, we sample $200$ random strings from each of the 22 sections of the Pile dataset [gao2020pile], and prompt GPT-4 to predict the source of the strings. Unlike the previous experiment, the ground-truth of string source is known in this validation experiment. Figure 5 illustrates the accuracy for each section, indicating that $50\%$ of the sections exhibit an accuracy rate of at least $90\%$ with $4.5\%$ being the base accuracy of a random predictor. Furthermore, GPT-4 predicts the correct source on an average of $78\%$ strings across all $22$ sections of the Pile dataset. In addition, misclassified strings are often assigned to sections of a similar category, e.g., NIH Explorer misclassified as PubMed Central (more details in Appendix 13). Therefore, the GPT-4 predictions can be considered as reliable.
@@
Figure 6: Most participants classified the memorized sequences detected by [biderman2024emergent] as non-sensitive, with fewer than 10% marking them as privacy-sensitive, indicating that the perceived privacy risk of these strings is generally low.
Finally, we conduct a human survey on Prolific to evaluate the extent of sensitive information present in randomly chosen $100$ memorized strings from [biderman2024emergent]. The survey results are summarized in Figure 6. The majority of participants classified memorized strings as non-sensitive, while $10\%$ participants disagree and mark the strings as containing privacy-sensitive information. Therefore, most crowdsourced participants do not perceive the sampled stings as containing privacy-sensitive content. Further details on the setup are provided in Appendix 15. Thus, by distinguishing between sensitive and non-sensitive entities, we demonstrate a deeper understanding of actual privacy threat.
4 Privacy-Utility-Efficiency Interplay
@@
We use the distinction between sensitive and non-sensitive tokens to study the privacy and utility impact of training models with three different fine-tuning methods: full fine-tuning (FFT), Differentially Privacy (DP-SGD), and Low-Rank Adaptation (LoRA). We also investigate the computational efficiency of each method. Our goal is to answer the following questions: “How prone is each method to recollecting the sensitive parts of the training data? (Privacy)” , “How effective is each method at predicting non-sensitive parts of test data? (Utility)” , “What is the computational cost associated with each method? (Efficiency)”.
@@
To answer these questions, we use each fine-tuning method to train the four models – Pythia-1B [biderman2023pythia], Gemma-2B [team2024gemma], Llama2-7B [touvron2023llama], and Qwen2.5-7B [yang2024qwen2] on two datasets, CustomerSim and SynBio. More information about the datasets and our methodology for distinguishing between sensitive and non-sensitive tokens can be found in Section 3.2. We train the models for 50 epochs on each dataset. Details on hyperparameters can be found in Appendix 12.
@@
For each fine-tuning method, model and dataset we report three metrics: 1) privacy as the loss on sensitive tokens (annotated by GPT-4) in the training data, 2) utility as the loss on non-sensitive tokens (the remaining tokens) on a held-out test set, and 3) efficiency based on the relative amount of computation and memory usage of each method. Additionally, to assess how fine-tuning affects the underlying abilities and knowledge of the base model, we measure the performance of the fine-tuned Gemma model (trained with CustomerSim data) on general language understanding benchmarks: SCIQ [SciQ], a dataset of over 13,000 crowdsourced questions on Physics, Chemistry, and Biology; MMLU [hendryckstest2021, hendrycks2021ethics], a large multi-task dataset covering various domains of knowledge; and Hellaswag [zellers-etal-2019-hellaswag], a dataset for commonsense inference.
[…]
@@
Efficiency: The efficiency of each method is determined by the amount of computation it requires, and also other factors such as memory requirements, which can affect the usable batch-size and thus the overall training throughput. Following [kaplan2020scaling], we estimate the amount of training compute ($C$) in floating point operations (FLOPs) for full fine-tuning as $C_{\text{FFT}}=6DN$, where $D$ is the number of training tokens and $N$, the number of model parameters. For each method, we report its compute requirements relative to the FFT-baseline based on measurements using the PyTorch profiler [^2] https://pytorch.org/docs/stable/profiler. We also comment on other factors that affect training throughput.
4.1 Full fine-tuning for utility costs privacy
@@
Efficiency: FFT serves as our efficiency baseline. It has moderate compute requirements (discussed above), and relatively high memory requirements, since in addition to the input-dependent activations, we need to keep four numbers per model parameter in GPU memory: the parameter value, its gradient, and two optimizer states (first and second moments of the gradient for Adam [kingma2014adam]).
@@
(a) Pythia (CustSim)
(b) Gemma (CustSim)
(c) Llama2 (CustSim)
(d) Qwen2.5 (CustSim)
(e) Pythia (SynBio)
(f) Gemma (SynBio)
(g) Llama2 (SynBio)
(h) Qwen2.5 (SynBio)
Figure 9: Privacy–utility trade-offs for differential privacy (DP-SGD) across models and datasets with varying noise levels ($\sigma$). Larger $\sigma$ increases privacy but reduces utility, with the decline more pronounced in larger models. On SynBio, DP-SGD exhibits a similar pattern. Gemma shows a unique utility drop after two epochs, highlighting the complexity of the privacy-utility trade-off.
4.2 DP-SGD fine-tuning for privacy and utility costs efficiency
Differential Privacy (DP-SGD) algorithms  [10.1145/2976749.2978318] aim to safeguard the privacy of individual training data points by limiting their influence on the gradient updates during training.
Update rules: DP-SGD clips the $l_{2}$ norm of each data point’s gradient at a threshold $T$, followed by adding noise with magnitude $\sigma$ to each clipped gradient. The purpose of clipping is to reduce data sensitivity by ensuring that the impact of data points with high gradient magnitudes on the model parameters is limited. Adding noise further obscures the contribution of individual data points, making it difficult to infer specific data points from the model.
@@ 4.2 DP-SGD fine-tuning for privacy and utility costs efficiency
We vary the noise hyperparameter $\sigma$ for the experiments. The clipping gradient norm $T$ is fixed at $10^{-2}$, as in [shi-etal-2022-selective].
@@
Figure 10: Fine-tuning Gemma with DP-SGD results in a substantial accuracy declines of 70%, 10%, and 30% across the benchmarks.
Privacy-Utility trade-off: Figures 9(a), 9(b), and 9(c) illustrate the privacy-utility trade-off when varying the noise $\sigma$ on the CustomerSim dataset across the Pythia, Gemma, and Llama2 models. As in previous plots, each curve begins with the performance of the pre-trained model. It is evident that DP-SGD maintains privacy effectively with minimal degradation. But there is a trade-off between privacy and utility. For all models, lower noise values such as $\sigma=0.1$ are able to achieve better utility than higher ones such as $\sigma=0.5$ and $\sigma=0.9$, but they also decrease privacy more. While the total amount of utility achievable with DP-SGD is limited, especially for the larger Llama2-7B model, overall, it provides good privacy-utility trade-offs. Similar trends can be observed for the SynBio dataset in Figures 9(e) and 9(g), but for Figure 9(f) (with Gemma) utility declines after approximately two epochs.
Impact on benchmark datasets: Figures 10(a), 10(b) and 10(c) show the benchmark accuracy over epochs for the DP-SGD fine-tuned Gemma model with $\sigma=0.1$ on SCIQ, MMLU, and HellaSwag datasets, respectively. Accuracy drops gradually with increasing epochs for all benchmarks, stabilizing at lower levels, indicating that fine-tuning with DP-SGD significantly reduces the knowledge retention capacity.
Efficiency: Differential privacy comes with a high computational cost, since the gradients of each datapoint need norm and clipping computations, and additional noise values are added. Empirically, we observe a relative FLOPs requirement of $C_{\text{DP-SGD}}/C_{\text{FFT}}=1.33$. Additionally, the per-sample operations required for clipping mean that we need to keep copies of the gradient for each datapoint in memory, which requires substantially more GPU memory, which decreases the feasible batch size and the overall training throughput even further.
Takeaway: DP-SGD offers a reasonable privacy-utility trade-off and is less susceptible to learning sensitive data. Increasing noise rates lead to poor utility across all models. Convergence with DP-SGD is not guaranteed, particularly in larger models such as Llama2. The privacy gains come at the cost of efficiency. Additionally, fine-tuning with DP-SGD leads to quick decline in the benchmark performance. The degree of measures along the tradeoff, knowledge retention, and efficiency are:
Utility-privacy trade-offs: moderate
Retention of base performance: poor
Efficiency: poor
(a) Pythia (CustSim)
(b) Gemma (CustSim)
(c) Llama2 (CustSim)
(d) Qwen2.5 (CustSim)
(e) Pythia (Rank 32)
(f) Gemma (Rank 32)
(g) Llama2 (Rank 32)
(h) Qwen2.5 (Rank 32)
Figure 11: Privacy–utility trade-offs for LoRA fine-tuning with ranks 16 and 32 on CustomerSim and varying scaling factor $\alpha$. Increasing $\alpha$ generally improves or maintains utility but reduces privacy. Smaller models achieve more favorable privacy–utility trade-offs, while larger models retain utility at the cost of reduced privacy during extended training.
@@
(d) Qwen2.5 (Rank 16)
(e) Pythia (Rank 32)
(f) Gemma (Rank 32)
(g) Llama2 (Rank 32)
(h) Qwen2.5 (Rank 32)
Figure 12: Privacy–utility trade-offs for LoRA fine-tuning with ranks 16 and 32 on SynBio and varying scaling factor $\alpha$. Increasing $\alpha$ generally improves or maintains utility but reduces privacy. Smaller models achieve more favorable privacy–utility trade-offs, while larger models retain utility at the cost of reduced privacy during extended training.
@@
4.3 LoRA fine-tuning for efficiency and utility improves privacy
LoRA [hu2022lora] is a parameter-efficient fine-tuning method developed to reduce the compute and memory requirements of fine-tuning LLMs, and to reduce the size of storing fine-tuned checkpoints. It enjoys large popularity for LLM fine-tuning.
@@ 4.3 LoRA fine-tuning for efficiency and utility improves privacy
We hypothesize that LoRA’s low-rank updates restrict the model’s capacity to memorize precise details, which could have an effect similar to the noisy updates in DP-SGD. Therefore, LoRA has the potential to provide better privacy-utility trade-offs than FFT, similar to DP-SGD, while also being computationally more efficient.
Privacy-Utility trade-off: We investigate the effects of varying both the rank $r$ and scaling parameter $\alpha$ of LoRA. We use common rank values of $16$ and $32$ and $\alpha\in\{16,32,64,128\}$. While LoRA has been explored for privacy in conjunction with DP-SGD [yu2022differentiallyprivatefinetuninglanguage], there has been no prior work that specifically examines the privacy benefits of LoRA alone.
Figures 11(a)11(c) show the trade-off with rank $16$ and varying parameters of $\alpha$ for the CustomerSim dataset across Pythia, Gemma and Llama2 models. As seen in Figures 11(a) and 11(b), smaller-scale models (Pythia and Gemma) exhibit a better privacy-utility trade-off when the rank and $\alpha$ values are equal, compared to the other configurations. For the larger model, the privacy declines at later epochs, while utility is mostly retained. This trend is similar to the one observed for rank $32$ in Figures 11(e)11(g). We also analyze the trade-off for the SynBio dataset in Figures 12(a)12(d) for rank 16, and Figures 12(e)12(g) for rank 32, which make similar observations. However, due to the more unstructured nature of the SynBio dataset, there is a larger reduction in utility after certain epochs compared to the CustomerSim dataset.
@@
Efficiency: During training, LoRA has slightly larger compute requirements for the forward pass than full fine-tuning, since additional FLOPs are required for the adapter matrices $\Delta W$, though they are much smaller than the full base weight $W_{0}$. However, during the backward pass, LoRA requires less compute, since no gradients need to be computed for the base weights $W_{0}$. We observe a relative FLOPs requirement of $C_{\text{LoRA}}/C_{\text{FFT}}\approx 0.65$. The original paper [hu2022lora] reports a 25% speedup during training. LoRA has needs of less GPU memory than FFT, since no optimizer states and gradients need to be stored for the base weights $W_{0}$, which makes it possible to run it with larger batch-sizes and thus an overall increased training throughput.
Takeaway: We are one of the first works to explore the privacy benefits of parameter-efficient fine-tuning methods, particularly LoRA. We vary LoRA’s $r$ and $\alpha$ hyperparameters and observe that all configurations are able to achieve high utility, while especially lower $r$ and $\alpha$ values also preserve a high degree of privacy. For each model, the optimal privacy-utility trade-off value is achieved with $r=\alpha$. In addition, we observe that LoRA, after being fine-tuned on the CustomerSim dataset, did not lose much of its abilities on the benchmark datasets, maintaining results comparable to those of pre-trained model. Finally, LoRA is much more computationally and memory efficient than FFT and especially DP-SGD. Overall it provides the best trade-offs in terms of utility, privacy and efficiency and shows that privacy can be achieved without additional computational costs. The degree of measures along the trade-off, knowledge retention, and efficiency are: